OpenStudy (anonymous):

Which of the following equations have the same solution? Solve all three equations separately and explain each step. I. 2x + 4 = 3x – 11 II. 3x + 3 = 51 III. the quantity of 2x plus 1, divided by 3 equals the quantity of x plus 6, divided by 2

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

solve for the x for each one of them

OpenStudy (anonymous):

2x+4 = 3x-11 combine like terms 15 = x -- answer 3x+3=51 --subtract 3 3x = 48 -- divide 3 x=16 -- answer 2x+1/3=x+6/2 -- cross multiply 2(2x+1) = (x+6)3 -- distribute the 3 and 2 4x+2 = 3x+18 -- combine like terms x = 16 -- answer